Lactic acid, the substance that causes sour milk to taste as it does, has a formula weight of 74. It consists of 48.63% C and 8.18% H, with the rest being oxygen. Calculate the empirical and molecular formulas of lactic acid, arranging the atomic symbols in alphabetical order.
1
Expert's answer
2019-06-17T02:54:10-0400

"\% O= 100 \% - (48.63 \% + 8.18\%)=43.19 \%"

"\dfrac{74g/mol*48.63\%}{12g/mol*100\%}=2.998(~3)atoms(C)"

"\dfrac {74g/mol*8.18\%}{1g/mol*100\%}=6.05 atoms(H)"

"\dfrac{74g/mol*43.19 \%}{16g/mol*100\%}" ~ 3 atoms(O)

molecular formula of lactic acid: C3H6O3